OpenStudy (jennychan12):

Find the exact coordinates of the centroid for the region bounded by the curves y=x+2 and y=x^2.

OpenStudy (jennychan12):

I already found the x coordinate, but how do you find the y coordinate?

OpenStudy (jennychan12):

It's from pretty little liars.

OpenStudy (abb0t):

y =x+2 is above y = x\(^2\) in the interval [-1, 2] (you can sketch the graph and check those points). Centroid = (xbar, ybar) where xbar =\(\sf \color{}{\frac{M_y}{A}}\) and ybar = \(\sf \color{}{\frac{M_x}{A}}\) With M\(_y\) = [ x (f(x) - g(x)) dx ] from -1<x<2 and same for M\(_x\) = ∫[ \(\sf \color{}{\frac{1}{2}}\) ((f(x))\(^2\) - (g(x))\(^2\)dx on -1<x<2

OpenStudy (jennychan12):

i already found xbar = \[\frac{ \int\limits\limits_{-1}^{-2}x(x+2-x^2)dx }{ \int\limits\limits_{-1}^{2}(x+2-x^2)dx }\] = 0.5 :) i just need to find ybar, which i'm not really sure how to find :(